# BitDelta: Your Fine-Tune May Only Be Worth One Bit

### v2

@@ Abstract
Large Language Models (LLMs) are typically trained in two phases: pre-training on large internet-scale datasets, and fine-tuning for downstream tasks. Given the higher computational demand of pre-training, it’s intuitive to assume that fine-tuning adds less new information to the model, and is thus more compressible. We explore this assumption by decomposing the weights of fine-tuned models into their pre-trained components and an additional delta. We introduce a simple method, BitDelta, which successfully quantizes this delta down to 1 bit without compromising performance. This interesting finding not only highlights the potential redundancy of information added during fine-tuning, but also has significant implications for the multi-tenant serving and multi-tenant storage of fine-tuned models. By enabling the use of a single high-precision base model accompanied by multiple 1-bit deltas, BitDelta dramatically reduces GPU memory requirements by more than 10$\times$, which can also be translated to enhanced generation latency in multi-tenant settings. We validate BitDelta through experiments across Llama-2 and Mistral model families, and on models up to 70B parameters, showcasing minimal performance degradation over all tested settings. Code is avaliable at https://github.com/FasterDecoding/BitDelta.
Machine Learning, ICML
Machine Learning, ICML
@@ 1 Introduction
After large-scale pretraining, foundation models are typically fine-tuned for specific downstream tasks (Devlin et al., 2019; Radford et al., 2018, 2019). This pretrain-finetune paradigm has revolutionized machine learning; LLMs have not only proven effective for critical tasks such as instruction following and alignment (Ouyang et al., 2022), but are also performant on a wide array of niche yet highly impactful applications (Xu et al., 2024; Qiu et al., 2023). Through fine-tuning, LLMs are adeptly equipped to align with distinct user preferences or specialized task requirements, showcasing an unprecedented level of adaptability. Thus, the prospect of serving millions of uniquely fine-tuned models, each tailored to individual tasks and user needs, presents a promising vision for the future of machine learning.
@@
From the delta decomposition point of view, parameter-efficient fine-tuning (PEFT) methods like LoRA (Hu et al., 2021; Houlsby et al., 2019a; Rebuffi et al., 2017; Dettmers et al., 2023; Chen et al., 2023d) effectively enforce a highly structured and compressed form of delta during fine-tuning, a powerful insight for model serving of PEFT-based fine-tunes. (Sheng et al., 2023) and (Chen et al., 2023b) explore multi-tenant serving of LoRA-based fine-tunes.
@@
Nevertheless, recent work has shown that PEFT methods may not yet match the model quality of full parameter fine-tuning, especially on high resource tasks (Chen et al., 2022), and are fairly sensitive to hyperparameter choice and prompting methods (Niederfahrenhorst et al., 2023). As a result, we notice that among the 2307 LLMs (as of time of writing) on the Open LLM Leaderboard (Beeching et al., 2023) with a valid README file, only $<20\%$ indicate that they exclusively use LoRA. Most models are full parameter fine-tunes, model merges (Yu et al., 2023; Jin et al., 2023; Wortsman et al., 2022) of full parameter fine-tunes, or model merges of LoRA based fine-tunes (which are effectively high-rank).
It is also attractive to approximate general deltas with low-rank matrices post-training. However, experimental results show that this is challenging (Table 1), as deltas from full parameter fine-tunes tend to be fairly high-rank (Figure 2).
@@
Figure 2: CEV plot of a $4096\times 4096$ weight delta between Llama 2-7B and Vicuna-7B v1.5. Deltas from full parameter fine-tuning are fairly high rank, making low-rank approximations difficult.
@@
Instead, we draw from the insight that motivates PEFT methods in general: Given the higher computational demand of pre-training, it’s intuitive to assume that fine-tuning adds less new information to the model, and is thus much more compressible. Indeed, we find that we can efficiently quantize the delta to merely 1 bit with almost no performance drop. We propose BitDelta, an efficient post-training quantization (PTQ) solution that acts on the weight delta between a fine-tuned model and its underlying base model.
BitDelta has two stages: 1) We quantize the delta of each weight matrix into a scaling factor multiplied by a binary matrix. Specifically, we take the sign of the delta to form the binary matrix and initialize the scaling factor as the average of the absolute values of the delta, which minimizes $L_{2}$ norm. 2) We further calibrate the scaling factors through model distillation over a small calibration dataset while keeping the binary matrices frozen. Despite the small number of trainable parameters and training steps, we find that this distillation process is effective in further recovering model quality. BitDelta is extremely efficient compared to other quantization methods; we are able to compress 70B models in roughly 10 minutes, whereas methods like GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2023) may take multiple GPU hours. Our experiments over 17 popular fine-tuned models affirm that BitDelta can be applied across various model types and sizes with minimal impact on performance.
BitDelta opens up opportunities to efficiently serve multiple fine-tuned models with shared servers: By only storing a single full-precision base model, and (dynamically) loading and performing batched inference over multiple 1-bit deltas, we can efficiently represent multiple fine-tuned models. Compared to naively using full precision fine-tuned models, deltas compressed by BitDelta are more than 10$\times$ smaller, consuming less GPU memory, and can therefore be loaded faster. This addresses the storage challenge. Moreover, since LLM inference is memory-bound (Leviathan et al., 2022; Chen et al., 2023a; Cai et al., 2024), the latency of each decoding step is proportional to the GPU memory consumption of the model weights. With an efficient CUDA kernel implementation, we can translate this memory reduction into a latency speedup, similar to other quantization methods (Frantar et al., 2022; Lin et al., 2023). As a preliminary attempt, we implement a 1-bit matrix multiplication kernel in Triton (Tillet et al., 2019) and show that it can improve the multi-tenant serving latency by about 2$\times$, addressing the serving challenge.
Finally, we study a few extensions of BitDelta, where we quantize the base model, and where we iteratively apply BitDelta. Experimental results show that our method is quite general and can be applied to various use cases.
@@
Table 1: Comparison between BitDelta and a SVD based method ($r=16$), with Llama 2-7B and Vicuna-7B v1.5 as the base and fine-tuned models. BitDelta is performant across the board, whereas the SVD-based method fails to sufficiently capture the fine-tuned information.
Model/Method Train Loss TruthfulQA GSM8K MT-Bench Adjusted Average2.1 $\uparrow$
@@
Vicuna-7B v1.5 50.36 19.03 6.04 60.51
BitDelta-Initial 0.41 47.63 19.56 5.67 60.99
BitDelta 0.052 49.97 20.17 5.99 60.68
SVD-Initial 0.78 44.56 15.92 4.45 60.60
SVD 0.34 44.55 14.18 4.98 60.20
@@ 2.1 Full Model Compression
Quantization techniques are widely used to reduce memory consumption and improve LLMs’ generation latency. Xiao et al. (2023) implement a technique that rescales between activations and parameters, effectively mitigating outlier activations to facilitate smoother quantization. Dettmers et al. (2022) develop an approach that decomposes matrix multiplications into 8-bit computations, with an additional 16-bit process for handling outliers. Exploring further, Frantar et al. (2022) introduce a method that iteratively rounds weight columns to 3-4 bits of precision. Similarly, Lin et al. (2023) propose an activation-aware quantization scheme that selectively preserves crucial weights while compressing the majority to 3-4 bits. In a different vein, Kim et al. (2023) devise a sparse, low-precision pattern focusing on a small yet significant set of weights. Lastly, Chee et al. (2023) utilize incoherence processing to quantize model weights to as low as 2 bits with minimal impact on performance.
@@
Pruning also aims to reduce the memory consumption of neural networks. It accomplishes this by pushing certain parameter values to zero, inducing sparsity in the model (LeCun et al., 1989; Han et al., 2015, 2016; Zhu & Gupta, 2017). However, these methods may fail to take advantage of modern hardware like GPUs unless using certain structured sparsity patterns like 2:4 (50%) sparsity (Mishra et al., 2021). Frantar & Alistarh (2023) demonstrate a pruning method on LLMs that successfully utilizes the 2:4 sparsity pattern and achieves a 50% sparsity ratio. It is challenging to obtain higher sparsity while being hardware-friendly.
2.2 Delta Compression
Parameter-efficient fine-tuning.
Parameter-efficient fine-tuning (PEFT) techniques reduce the number of trainable parameters during fine-tuning, reducing memory and compute demand while achieving promising accuracy. Houlsby et al. (2019b) introduce adapter layers, freezing the base model during fine-tuning. Hu et al. (2021) propose Low-Rank Adaptation (LoRA), which enforces that weight updates during fine-tuning are low rank.
By enforcing structure in the fine-tuning information, PEFT techniques can viewed as compression-aware training. Similar to our method, such compression induced by PEFT techniques can also enhance multi-tenant serving. Chen et al. (2023b) and Sheng et al. (2023) explore this idea by developing scalable multi-tenant systems for LoRA-based fine-tunes, based on customized CUDA kernels that exploit the low-rank structure of the fine-tuned weight deltas.
@@ 2.2 Delta Compression
Most related to our work, a few studies explore the idea of post-training delta compression by adopting existing compression techniques like GPTQ, unstructured pruning (Han et al., 2016), or even classic lossless compression algorithms. Isik et al. (2023) focus on reducing the delta size to save storage. Yu et al. (2023) utilize pruning to improve model merging applications. Yadav et al. (2023) reduces the size of PEFT modules to save storage. Ryu et al. (2023) combines quantization with low-rank approximation to reduce the delta size. The concurrent and independent work by Yao & Klimovic (2023) also explores using delta compression to improve multi-tenant serving, but focuses more on reducing the model loading time from disk to GPU. Compared to existing work, we offer a much simpler and faster method, BitDelta, achieving a more than 10$\times$ compression ratio while also being friendly to modern accelerators.
@@ 3.1 Method
BitDelta consists of two stages: 1) We quantize each weight matrix into a scalar multiplied by a binary matrix***In our experiments, we only quantize the linear layers in the Transformer blocks as they contribute most of the parameters and computation.. 2) We further calibrate the scalar factors with distillation. We describe each stage in this section:
@@
Let $W_{\text{base}},W_{\text{fine}}\in\mathbb{R}^{n\times m}$ be weight matrices from the base model and fine-tuned model, respectively. We define the weight delta as $\Delta=W_{\text{fine}}-W_{\text{base}}$, representing the modification in weights post-fine-tuning. For efficient representation, we aim to obtain a binarized estimator of this weight delta, denoted as $\hat{\Delta}$, by encoding its sign bits:
@@
and $\alpha$ is a high-precision scaling factor for the entire matrix. To minimize the quantization error in $L_{2}$ norm:
$\left\|\Delta-\hat{\Delta}\right\|_{2}^{2}=\sum_{ij}\left(\left|W_{ij}\right|-\alpha\right)^{2},$ (3)
we take
$\alpha=\frac{1}{nm}\sum_{ij}|W_{ij}|.$ (4)
@@
Table 2: BitDelta works on Llama-2 and Mistral families and on a wide range of model sizes ranging from 7B to 70B parameters. BitDelta works for many types of fine-tuned information, including SFT-based methods, RLHF-based methods, and context extension methods (RoPE scaling). Scale distillation is effective; it raises TruthfulQA/GSM8K scores to within 1-2 points of the baseline fine-tune, and MT-Bench scores to within 0.1-0.2 points.
Model Method TruthfulQA GSM8K MT-Bench Adjusted Average2.1 $\uparrow$

### v3

@@ Abstract
Large Language Models (LLMs) are typically trained in two phases: pre-training on large internet-scale datasets, and fine-tuning for downstream tasks. Given the higher computational demand of pre-training, it is intuitive to assume that fine-tuning adds less new information to the model, and is thus more compressible. We explore this assumption by decomposing the weights of fine-tuned models into their pre-trained components and an additional delta. We introduce a simple post-fine-tuning method, BitDelta, which successfully quantizes this delta down to 1 bit without compromising performance. This interesting finding not only highlights the potential redundancy of information added during fine-tuning, but also has significant implications for the multi-tenant serving and multi-tenant storage of fine-tuned models. By enabling the use of a single high-precision base model accompanied by multiple 1-bit deltas, BitDelta dramatically reduces GPU memory requirements by more than 10$\times$, thus reducing per-user generation latency by more than 10×10$\times$10 × in multi-tenant settings. We validate BitDelta through experiments across Llama-2, Mistral and MPT model families, and on models up to 70B parameters, showcasing minimal performance degradation in all tested settings.
@@ 1 Introduction
After large-scale pretraining, foundation models are typically fine-tuned for specific downstream tasks [16, 43, 44]. This pretrain-finetune paradigm has revolutionized machine learning; LLMs have not only proven effective for critical tasks such as instruction following and alignment [39], but are also performant on a wide array of niche yet highly impactful applications [61, 42]. Through fine-tuning, LLMs are adeptly equipped to align with distinct user preferences or specialized task requirements, showcasing an unprecedented level of adaptability. Thus, the prospect of serving millions of uniquely fine-tuned models, each tailored to individual tasks and user needs, presents a promising vision for the future of machine learning.
@@
From the delta decomposition point of view, parameter-efficient fine-tuning (PEFT) methods like LoRA [25, 24, 46, 15, 9] effectively enforce a highly structured and compressed form of delta during fine-tuning, a powerful insight for model serving of PEFT-based fine-tunes. Sheng et al. [49] and Chen et al. [7] explore multi-tenant serving of LoRA-based fine-tunes.
@@
Nevertheless, recent work has shown that PEFT methods may not yet match the model quality of full parameter fine-tuning, especially on high resource tasks [6], and are fairly sensitive to hyperparameter choice and prompting methods [38]. Biderman et al. [2] show that LoRA’s reduced expressivity, although providing desirable regularization, leads to significantly worse performance compared to full fine-tuning in math and programming tasks. As a result, we notice that among the 2307 LLMs (as of time of writing) on the Open LLM Leaderboard [1] with a valid README file, only $<20\%$ indicate that they exclusively use LoRA. Most models are full parameter fine-tunes, model merges [64, 28, 59] of full parameter fine-tunes, or model merges of LoRA based fine-tunes (which are effectively high-rank).
@@
Figure 2: Cumulative Explained Variance (CEV) plot of a $4096\times 4096$ weight delta between Llama 2-7B and Vicuna-7B v1.5. Deltas from full parameter fine-tuning are fairly high rank, making low-rank approximations difficult.
It is also attractive to approximate general deltas with low-rank matrices post-training (in particular, post-fine-tuning). However, experimental results show that this is challenging (Table 1), as deltas from full parameter fine-tunes tend to be fairly high-rank (Figure 2).
@@
We instead draw from the insight that motivates PEFT methods in general: Given the higher computational demand of pre-training, it is intuitive to assume that fine-tuning adds less new information to the model, and is thus much more compressible. In fact, we find that we can efficiently quantize the delta to merely 1 bit with almost no performance drop. We propose BitDelta, an efficient post-training quantization (PTQ) solution that acts on the weight delta between a fine-tuned model and its underlying base model.
@@
Table 1: Comparison between BitDelta and a SVD based method, with Llama 2-7B and Llama 2-7B Chat as the base and fine-tuned models. BitDelta is performant across the board, whereas the SVD-based method fails to sufficiently capture the fine-tuned information.
Model/Method TruthfulQA GSM8K MT-Bench Adjusted Average1 $\uparrow$
@@
Llama 2-7B Chat 45.32 22.74 6.56 59.81
BitDelta-Initial 41.10 18.27 6.31 60.70
BitDelta 44.95 20.24 6.47 59.88
SVD-Initial ($r=16$) 42.57 7.13 4.73 60.58
SVD ($r=16$) 42.42 5.05 4.99 60.71
SVD-Initial ($r=128$) 43.90 17.82 5.68 60.21
SVD ($r=128$) 43.32 11.83 5.85 60.58
BitDelta consists of two stages: 1) We quantize the delta between a fine-tuned model’s weight matrix and base model’s weight matrix into a scaling factor multiplied by a binary matrix. Specifically, we take the sign of the weight delta to form the binary matrix and initialize the scaling factor as the average of the absolute values of the delta, minimizing $L_{2}$ quantization error. 2) We further calibrate the scaling factors through model distillation over a small calibration dataset while keeping the binary matrices frozen. Despite the small number of trainable parameters and calibration steps, we find that this distillation process is effective in further recovering model quality. Our experiments over 17 popular fine-tuned models affirm that BitDelta can be applied across various model types and model sizes with minimal impact on performance.
BitDelta creates opportunities to efficiently serve multiple fine-tuned models with shared servers: By only storing a single full-precision base model, and (dynamically) loading and performing batched inference over multiple 1-bit deltas, we can efficiently represent multiple fine-tuned models. Compared to naively using full precision fine-tuned models, deltas compressed by BitDelta are more than 10$\times$ smaller, and can therefore be loaded faster. This addresses the storage challenge. Moreover, since LLM inference is memory-bound [32, 5, 3], the latency of each decoding step is proportional to the GPU memory consumption of the model weights. With an efficient CUDA kernel implementation, we can translate this memory reduction into a latency reduction, similar to other quantization methods [19, 33]. Using the $W_{INT1}A_{FP16}$ kernel from BitBLAS [58], we improve the multi-tenant serving latency of full-parameter fine-tuned models by more than 10×10$\times$10 ×.
Finally, we study a few extensions of BitDelta, where we quantize the base model and where we iteratively apply BitDelta. Experimental results show that our method is quite general and can be applied to various use cases.
@@ 2.1 Full Model Compression
Quantization techniques are widely used to reduce memory consumption and improve LLMs’ generation latency. Xiao et al. [60] implement a technique that rescales between activations and parameters, effectively mitigating outlier activations to facilitate smoother quantization. Dettmers et al. [14] develop an approach that decomposes matrix multiplications into 8-bit computations, with an additional 16-bit process for handling outliers. Exploring further, Frantar et al. [19] introduce a method that iteratively rounds weight columns to 3-4 bits of precision. Similarly, Lin et al. [33] propose an activation-aware quantization scheme that selectively preserves crucial weights while compressing the majority to 3-4 bits. Kim et al. [29] devise a sparse, low-precision pattern focusing on a small yet significant set of weights. Chee et al. [4] utilize incoherence processing to quantize model weights to as low as 2 bits with minimal impact on performance.
@@
Pruning also aims to reduce the memory consumption of neural networks. It accomplishes this by pushing certain parameter values to zero, inducing sparsity in the model [31, 21, 22, 67]. However, these methods may fail to take advantage of modern hardware like GPUs unless using certain structured sparsity patterns like 2:4 (50%) sparsity [36]. Frantar and Alistarh [18] demonstrate a pruning method on LLMs that successfully utilizes the 2:4 sparsity pattern and achieves a 50% sparsity ratio. It is challenging to obtain higher sparsity while being hardware-friendly.
@@
Most related to our work, a few studies explore the idea of post-training delta compression by adopting existing compression techniques like GPTQ, unstructured pruning [22], or even classic lossless compression algorithms. Isik et al. [26] focus on reducing the delta size to save storage. Yu et al. [64] utilize pruning to improve model merging applications. Yadav et al. [62] reduces the size of PEFT modules to save storage. Ryu et al. [47] combines quantization with a low-rank approximation to reduce the delta size. The concurrent and independent work by Yao and Klimovic [63] also explores using delta compression to improve multi-tenant serving, but focuses more on reducing the model loading time from disk to GPU. Compared to existing work, we offer a much simpler and faster method, BitDelta, achieving a compression ratio of more than 10$\times$ while also being friendly to modern accelerators.
@@ 3.1 Method
BitDelta consists of two stages: 1) We quantize each weight matrix into a scalar multiplied by a binary matrix***In our experiments, we only quantize the linear layers in the Transformer blocks as they contribute the majority of the parameters and computation.. 2) We further calibrate the scalar factors using model distillation. We describe each stage in this section:
@@
Let $W_{\text{base}},W_{\text{fine}}\in\mathbb{R}^{n\times m}$ be weight matrices from the base model and fine-tuned model respectively. We define the weight delta as $\Delta=W_{\text{fine}}-W_{\text{base}}$, representing the modification in weights post-fine-tuning. For efficient representation of this weight delta, we aim to obtain a binarized estimator by encoding its sign bits, denoted as $\hat{\Delta}$:
@@
and $\alpha$ is a high-precision scaling factor for the entire matrix. To minimize the quantization error of $\Delta$ in $L_{2}$ norm:
$\left\lVert\Delta-\hat{\Delta}\right\rVert_{2}^{2}=\sum_{ij}(|W_{ij}|-\alpha)^{2},$ (3)
we initialize $\alpha$ as follows:
$\alpha=\frac{1}{nm}\sum_{ij}|\Delta_{ij}|.$ (4)
@@
Table 2: BitDelta works on Llama-2 and Mistral families and on a wide range of model sizes ranging from 7B to 70B parameters. BitDelta works for many types of fine-tuned information, including SFT-based methods, RLHF-based methods, and context extension methods (RoPE scaling). Scale distillation is effective, raising TruthfulQA/GSM8K scores to within 1-2 points of the baseline fine-tune, and MT-Bench scores to within 0.1-0.2 points.
Model Method TruthfulQA GSM8K MT-Bench Adjusted Average1 $\uparrow$
@@
Scale distillation.
The scaling factor $\alpha$ intuitively plays a more significant role in the low-bit regime. Additionally, per-matrix $L_{2}$ weight error is not a perfect measure of degradation in overall model quality. We further optimize these scales by performing model distillation to align the output logits of the quantized model to that of the original fine-tuned model. More concretely, we freeze the model weights and optimize for the following objective:
$\boldsymbol{\alpha}^{*}=\arg\min_{\boldsymbol{\alpha}}\mathbb{E}_{x\sim\mathbf{X}}\left[\left\|\mathbf{Z}_{\text{fine}}(x)-\mathbf{Z}_{\text{bin}}(x;\boldsymbol{\alpha})\right\|^{2}\right]$ (5)
where $\mathbf{X}$ is a calibration dataset, and $\mathbf{Z}(\cdot)$ are the logits of the respective models. Scale distillation is fairly robust to choice $\mathbf{X}$, as 1) the process is extremely parameter efficient, and 2) the crucial aspect of the process is to logit match with the fine-tuned model, regardless of the actual text content.
For our experiments, we distill on the C4 dataset [45], consisting of generic internet data, using 800 samples of length 128. We use the same subset of C4 over all models to control for seed-based variations. We use the Adam optimizer [30] with $lr=10^{-4}$, $\beta=(0.9,0.999)$, $\epsilon=10^{-8}$. 1x80 GB A100 GPU is used to distill 7B and 13B models, and 6x80GB A100 GPUs are used to distill 70B models (2x for finetune, 4x for binarized). Scale distillation is fast; we can compress 70B models in roughly 10 minutes.
3.2 Methodology Cost
Compared to full parameter and parameter efficient fine-tuning methods, BitDelta is extremely cheap. While fine-tuning methods require training thousands to millions of parameters, BitDelta only necessitates training a single parameter per weight matrix. Moreover, BitDelta operates efficiently with input sequences of length 128, unlike fine-tuning methods that demand longer sequences to saturate the context window (4k, 8k, etc.). Crucially, BitDelta requires only 200 training steps (assuming a batch size of 4), which is significantly less compared to the 10000-1000000 steps at higher batch sizes needed by fine-tuning methods. Thus, in terms of methodology cost, we liken BitDelta more to post-training quantization (PTQ) schemes like GPTQ [19] and AWQ [33], rather than full parameter or parameter efficient fine-tuning, while being faster than most PTQ schemes.
Table 3: Continuation of Table 2.
Model Method TruthfulQA GSM8K MT-Bench Adjusted Average1 $\uparrow$
@@ 3.2 Methodology Cost
MPT-7B 33.37 6.22 57.95
MPT 7B-Chat Baseline 40.22 7.96 5.00 56.5
BitDelta-Initial 38.96 10.01 4.39 57.11
BitDelta 39.87 8.11 4.94 56.52
3.3 Implication
The ability to compress the delta to merely 1-bit opens up multiple opportunities for improving efficiency, enabling more effective model storage [26] – where a single base model can be maintained alongside multiple compressed deltas – and facilitating model hot-swapping  [7, 49]. With hot-swapping, the base model remains in GPU memory, and compressed deltas are dynamically loaded in accordance to incoming requests. In both cases, the compression ratio can be directly translated into reductions in storage needs and loading times.
Moreover, BitDelta enables the possibility of a multi-tenant serving system like Punica [7] or S-LoRA [49] but for general fine-tuned models instead of just LoRA models. Concretely, we consider the scenario where multiple models fine-tuned from the same base model are served with the same server. This setting greatly exploits the GPU resource and saves each fine-tuned model’s inference cost when their traffic is low or unbalanced. With BitDelta, we can keep one high-precision base model with multiple compressed deltas in the GPU memory. Compared to directly serving multiple fine-tuned models, this approach greatly saves memory consumption.
Since LLM inference follows the memory-bound computation pattern where the generation latency is proportional to the GPU memory used by the model weights, the lower memory consumption also suggests the opportunity to improve the serving latency. For example, Punica and S-LoRA exploit LoRA’s structure and memory saving by computing the activation product between the shared base weight, and low-rank fine-tuned delta weights separately. Similarly, we decompose the forward pass of each linear layer as follows:
where $X_{i}$ and $X_{i}^{\prime}$ represent input and output features to the $i$-th fine-tuned model, and the base model weight and the 1-bit delta are computed separately. For a batch of requests, $W_{\text{base}}X_{i}$ can be computed with the classic batched GEMM kernel. We utilize the BitBLAS  [58] $W_{INT1}A_{FP16}$ kernel that allows us to calculate $\hat{\Delta}_{i}X$ in a batched setting while keeping the 1-bit deltas quantized until they are transferred to the GPU cache. This kernel fuses the dequantization operation with the GEMM calculation, reducing the data moving overhead by a large factor.
4 Experiments
4.1 Setup